# Surfaces and other Peano Continua with no Generic Chains

### v1

@@ 6.1. Robust cycles
Broadly speaking, our goal is to isolate a region of $X$ (a circular subpartition to be precise) inside which there is well-defined notion of winding number, while also making sure that we can refine and extend given paths in such region so as to be able to reach arbitrarily large positive and negative values of winding number. Once this is ensured, if $(\mathcal{V},v)$ is as in the statement of Theorem 5.8, we can then find a brick partition $\mathcal{W}\preceq\mathcal{V}$ and two paths $w_{0},w_{1}$ on $\mathcal{W}$ refining $v$ and such that the winding number of $w_{0}$ is very large, while the one of $w_{1}$ is negative and very large in absolute value. If $X$ had a generic chain, Theorem 5.8 would permit to amalgamate up to distance one $(\mathcal{W},w_{0})$ and $(\mathcal{W},w_{1})$, and this would lead to a contradiction by Lemma 3.4, since $w_{0}$ and $w_{1}$ were chosen to have very distant winding numbers. The idea of using the winding number to show the failure of such amalgamation goes back to the proof of [GTZ]*Theorem 1.2.
@@
[scale=1]img/robust
Figure 1. A robust cycle.
Once this is done, all that is left is to show that the spaces considered in Theorem 1.2 satisfy the necessary property to perform this construction (this is done in Theorems 6.6, 6.10 and 6.12). Such property is formalized in Definition 6.3 below. The intuitive and simplified idea is to require the existence of a partition $\mathcal{S}$ of $X$ with a robust cycle $C$: by this we mean a cycle such that, given any brick partition $\mathcal{V}\preceq\mathcal{S}$ and walk $v$ on $\mathcal{V}$, even in the situation where $v$ intersects and apparently cuts $C$, it is always possible to zoom in (that is, find a refinement) to see that $C$ is still intact and in fact $v$ can be prolonged with a circular path confined in $C$ in a strong sense.
@@
Let us say that a Peano continuum $X$ has a robust cycle if there are a brick partition $\mathcal{S}$ with a spaced circular path $C$ of length $\ell\geq 4$, and a brick partition $\mathcal{U}\preceq\mathcal{S}$ with a walk $u$ on $\mathcal{U}$ such that for any brick partition and walk $(\mathcal{V},v)\preceq(\mathcal{U},u)$ there are $(\mathcal{V}^{\prime},v^{\prime}{}^{\frown}c)\preceq(\mathcal{V},v)$ where $v^{\prime}{}^{\frown}c$ is a lasso path such that:
@@
The existence of a robust cycle gives us a brick partition $\mathcal{S}$, a spaced circular path $C$ of length $\ell\geq 4$, a brick partition $\mathcal{U}\preceq\mathcal{S}$, and and a walk $u$ on $\mathcal{U}$. Up to taking a refinement, by Lemma 5.3 we can suppose that
@@
Suppose we are given a brick partition and a walk $(\mathcal{V},v_{0})\preceq(\mathcal{U},u)$. By definition of robust cycle and by (Spacing out Paths). (Lemma 5.4) we can assume, up to a refinement, that $v_{0}$ is a spaced lasso path $v{}^{\frown}c_{+}$ such that:
@@
Let $c_{-}$ be the spaced circular path going in the opposite direction of $c_{+}$. Set $N\coloneqq\operatorname{wd}_{C}(\rho^{\mathcal{V}}_{\mathcal{S}}\cdot c_{+})$, so that $\operatorname{wd}_{C}(\rho^{\mathcal{V}}_{\mathcal{S}}\cdot c_{-})=-N$. Up to switching the role of $c_{+},c_{-}$, we might assume that $N\geq\ell>0$.
@@
Since $v{}^{\frown}\bar{w}_{0}$ and $v{}^{\frown}\bar{w}_{1}$ are uncrossed, by (Upgrading Walks to Paths). (Theorem 6.2) twice, we find a brick partition $\mathcal{W}\preceq\mathcal{V}$ and spaced paths $w_{0},w_{1}$ on $\mathcal{W}$ such that $\rho^{\mathcal{W}}_{\mathcal{V}}\cdot w_{i}$ monotonically refines $v{}^{\frown}\bar{w}_{i}$. We stress that $w_{0},w_{1}$ are paths, not lassos.
Using Lemma 3.2 and the definitions of $w_{0},w_{1}$, we obtain the following computations of winding numbers:
@@
We now apply Theorem 5.8, hence towards a contradiction suppose there are a brick partition $\mathcal{Z}$, two monotone epimorphisms $\alpha_{0},\alpha_{1}\colon\mathcal{Z}\to\mathcal{W}$, and a reduced walk $z$ on $\mathcal{Z}$ such that $\alpha_{i}\cdot z\sqsubseteq w_{i}$, for $i=0,1$, and
@@
Notice that $\bar{z}_{v}$ might be the empty walk (this would be the case if $C=\mathcal{S}$). In particular, $\alpha_{0}\cdot\bar{z}_{0}$ refines and is confined to a final segment of $w_{0}$ whose image under $\rho^{\mathcal{W}}_{\mathcal{V}}$ contains $\bar{w}_{0}$, and moreover $\alpha_{0}\cdot\bar{z}_{0}(-1)=w_{0}(-1)$, so by 6.4.1.(b0) and Lemma 3.3
@@
$d_{\mathcal{U}}(\rho^{\mathcal{W}}_{\mathcal{U}}\alpha_{0}\cdot\bar{z}_{0}(j),\rho^{\mathcal{W}}_{\mathcal{U}}\alpha_{1}\cdot\bar{z}_{0}(j+1))\leq\\ \leq d_{\mathcal{U}}(\rho^{\mathcal{W}}_{\mathcal{U}}\alpha_{0}\cdot\bar{z}_{0}(j),\rho^{\mathcal{W}}_{\mathcal{U}}\alpha_{0}\cdot\bar{z}_{0}(j+1))+d_{\mathcal{U}}(\rho^{\mathcal{W}}_{\mathcal{U}}\alpha_{0}\cdot\bar{z}_{0}(j+1),\rho^{\mathcal{W}}_{\mathcal{U}}\alpha_{1}\cdot\bar{z}_{0}(j+1))\leq 2,$
@@
If $\bar{z}_{v}$ is the empty walk then $\operatorname{wd}_{C}(\rho^{\mathcal{W}}_{\mathcal{S}}\alpha_{1}\cdot\bar{z})=\operatorname{wd}_{C}(\rho^{\mathcal{W}}_{\mathcal{S}}\alpha_{1}\cdot\bar{z}_{0})$, so we obtain a contradiction from (6.6) and (6.9). Otherwise $\alpha_{1}\cdot\bar{z}_{v}$ refines an initial segment of $w_{1}$ and $\rho^{\mathcal{W}}_{\mathcal{V}}\alpha_{1}(\bar{z}_{v}(-1))$ is contained in $v$, and not in $\bar{w}_{1}$: indeed by definition of $\bar{z}_{0}$ the vertex $\rho^{\mathcal{W}}_{\mathcal{U}}(\alpha_{0}(\bar{z}_{v}(-1)))$ must be $E$-adjacent to one in $(\rho^{\mathcal{U}}_{\mathcal{S}})^{-1}(C)$, and this in combination with (6.5) yields
@@
Combining (6.9) and (6.10) we conclude $\operatorname{wd}_{C}(\rho^{\mathcal{W}}_{\mathcal{S}}\alpha_{2}\cdot\bar{z})>M_{v}+M_{c}$, again a contradiction with (6.6). ∎
@@ 6.2. Locally non-planar open subset
Proposition 6.5 (Disjoint Path Property).
@@
For each $1<i<\operatorname{lh}(u)$, let $x_{i}$ be a point in $\operatorname{Cl}(u(i-1))\cap\operatorname{Cl}(u(i))\cap(u(i-1)\vee u(i))$, and, for each $1<j<\operatorname{lh}(u^{\prime})$, let $x^{\prime}_{j}$ be a point in $\operatorname{Cl}(u^{\prime}(j-1))\cap\operatorname{Cl}(u^{\prime}(j))\cap(u^{\prime}(j-1)\vee u^{\prime}(j))$. We can assume that all these points are distinct by 4.2.(2), since $u$ and $u^{\prime}$ are paths, so for each $i<\operatorname{lh}(u)$ there is at most one $j<\operatorname{lh}(u^{\prime})$ such that $u(i)=u^{\prime}(j)$, and vice versa.
@@
Let $\lambda,\lambda^{\prime}$ be the two disjoint arcs resulting from the concatenations of the $\gamma_{i}$’s and $\gamma^{\prime}_{j}$’s, respectively. Let $\tilde{\mathcal{V}}\preceq\mathcal{U}$ be a brick partition fine enough that $\operatorname{Star}(\lambda,\tilde{\mathcal{V}})\cap\operatorname{Star}(\lambda^{\prime},\tilde{\mathcal{V}})=\emptyset$. Set $P_{i}\coloneqq\operatorname{Star}(\lambda,\tilde{\mathcal{V}}(u(i))),P^{\prime}_{j}\coloneqq\operatorname{Star}(\lambda^{\prime},\tilde{\mathcal{V}}(u^{\prime}(j)))$, for each $i<\operatorname{lh}(u),j<\operatorname{lh}(u^{\prime})$, and let $\mathcal{V}$ be the amalgam of $\tilde{\mathcal{V}}$ refining $\mathcal{U}$ and containing $\bigvee P_{i}$ and $\bigvee P^{\prime}_{j}$, for all $i<\operatorname{lh}(u),j<\operatorname{lh}(u^{\prime})$. Then $v\coloneqq\langle P_{0},\dots,P_{\operatorname{lh}(u)-1}\rangle$ and $v^{\prime}\coloneqq\langle P^{\prime}_{0},\dots,P^{\prime}_{\operatorname{lh}(u^{\prime})-1}\rangle$ are paths on $\mathcal{V}$ satisfying the conclusions of the proposition.
@@
The (Disjoint Path Property). allows us to find a robust cycle in any locally non-planar open set.
@@
First let us show that it is sufficient to find $(\mathcal{V}^{\prime},v^{\prime})\preceq(\mathcal{V},v)$ with $\mathcal{V}$ a brick partition and $v^{\prime}$ a path, and a circular path $c$ disjoint form $v^{\prime}$ satisfying (6.1) and (6.2) in Definition 6.3. Indeed, in this case, by (Path Doubling). (Proposition 6.1) we can assume that, up to a refinement, there are two paths $v^{\prime}_{0},v^{\prime}_{1}$ which are disjoint except in their endpoints, disjoint from $c$, and such that $\rho^{\mathcal{V}^{\prime}}_{\mathcal{V}}\cdot v^{\prime}_{i}\sqsubseteq v$. Let $P$ be a path from $v^{\prime}_{0}\cup v^{\prime}_{1}$ to $c$, and say without loss of generality that it starts in $v^{\prime}_{0}$. Then the lasso path that follows $v^{\prime}_{1}$ from beginning to end, then $v^{\prime}_{0}$ backwards until meeting $P$, then follows $P$, then does a complete lap following $c$, is as required.
We now proceed to find such $v^{\prime}$ and $c$. Suppose that $O\subseteq X$ is open and locally non-planar. Fix any brick partition $\mathcal{W}$ of $X$ which is fine enough that there are $W,W^{\prime}$ which are in the same $E$-connected component of $\mathcal{W}(O)$ but are not adjacent. Let $w$ be a path from $W$ to $W^{\prime}$ in $\mathcal{W}(O)$. By (Path Doubling)., there are a brick partition $\mathcal{S}\preceq\mathcal{W}$ and two paths $s_{0},s_{1}$ on $\mathcal{S}$ which are disjoint except for their endpoints and such that $\rho^{\mathcal{S}}_{\mathcal{W}}\cdot s_{i}$ monotonically refines $w$, for $i<2$. Let $C$ be the circular path which follows $s_{0}$ forward and then $s_{1}$ backwards. Up to a refinement, we can assume $C$ is spaced, by (Spacing out Paths). (Lemma 5.4). Since $s_{0},s_{0}$ have length at least $3$, $C$ has length at least $4$.
@@
Let $u$ be the empty walk and suppose that a brick partition and a walk $(\mathcal{V},v)\preceq(\mathcal{U},u)$ are given. Up to a refinement, we can suppose that $v$ is a path, by (Upgrading Walks to Paths). (Theorem 6.2). Use Lemma 5.3 to find a circular path $D^{\prime}$ on $\mathcal{V}$ such that $\rho^{\mathcal{V}}_{\mathcal{U}}\cdot D^{\prime}$ monotonically refines $D$. Since $D^{\prime}$ is confined in $\mathcal{V}(O)$, by (Disjoint Path Property). (Proposition 6.5) there are a brick partition $\mathcal{V}^{\prime}\preceq\mathcal{V}$ and two disjoint paths $v^{\prime},c$ on $\mathcal{V}^{\prime}$ such that $\rho^{\mathcal{V}^{\prime}}_{\mathcal{V}}\cdot v_{0},\rho^{\mathcal{V}^{\prime}}_{\mathcal{V}}\cdot c$ monotonically refine $v$ and $D^{\prime}$, respectively, and $c$ is circular. Then $c$ satisfies (6.1) by (6.11), and it follows by Lemma 3.2 that it satisfies (6.2). ∎
6.3. Planar open set with non-locally separating simple closed curve
The next two results capture the essential aspects behind the assumption of the existence of a non-locally separating simple closed curve in Theorem 1.2. We start with an auxiliary definition.

### v3

@@ 6.2. Robust cycles
Broadly speaking, our goal is to isolate a region of $X$ (a circular subpartition to be precise) inside which there is a well-defined notion of winding number, while also making sure that we can refine and extend given paths in such region so as to be able to reach arbitrarily large positive and negative values of winding number. Once this is ensured, if $(\mathcal{V},v)$ is as in the statement of Theorem 5.9, we can then find a brick partition $\mathcal{W}\preceq\mathcal{V}$ and two paths $w_{0},w_{1}$ on $\mathcal{W}$ refining $v$ and such that the winding number of $w_{0}$ is very large, while the one of $w_{1}$ is negative and very large in absolute value. If $X$ had a generic chain, Theorem 5.9 would permit to amalgamate up to distance one $(\mathcal{W},w_{0})$ and $(\mathcal{W},w_{1})$, and this would lead to a contradiction by Lemma 3.5, since $w_{0}$ and $w_{1}$ were chosen to have very distant winding numbers. The idea of using the winding number to show the failure of such amalgamation goes back to the proof of [GTZ]*Theorem 1.2.
@@
[scale=1]img/robust-v3
Figure 1. A sketch of a robust cycle: only the relevant elements of the brick partitions appearing in Definition 6.3 are shown, and the lasso $v^{\prime}{}^{\frown}c$ is depicted as a non-self-intersecting curve.
Once this is done, all that is left is to show that the spaces considered in Theorem 1.2 satisfy the necessary property to perform this construction (this is done in Theorems 6.6, 6.10 and 6.12). Such property is formalized in Definition 6.3 below. The intuitive and simplified idea is to require the existence of a partition $\mathcal{S}$ of $X$ with a robust cycle $C$: by this we mean a cycle such that, given any brick partition $\mathcal{V}\preceq\mathcal{S}$ and walk $v$ on $\mathcal{V}$, even in the situation where $v$ intersects and apparently cuts $C$, it is always possible to zoom in (that is, find a refinement) to see that $C$ is still intact and in fact $v$ can be prolonged with a disjoint circular path refining and confined in $C$ in a strong sense.
Such robust cycles do not exist, for instance, in the $2$-sphere $S^{2}$. Indeed, any circular path $C$ on a brick partition of $S^{2}$ can be separated by an arc, and by all paths $v$ on sufficiently fine brick partitions containing such arc. The resulting cut cannot be mended: any circular path which is confined to and winds around $C$, no matter on how fine a brick partition, must intersect $v$.
@@
Let us say that a Peano continuum $X$ has a robust cycle if there are a brick partition $\mathcal{S}$ with a spaced circular path $C$ of circular length $\ell\geq 4$, and a brick partition $\mathcal{U}\preceq\mathcal{S}$ with a walk $u$ on $\mathcal{U}$ such that for any brick partition and walk $(\mathcal{V},v)\preceq(\mathcal{U},u)$ there are $(\mathcal{V}^{\prime},v^{\prime}{}^{\frown}c)\preceq(\mathcal{V},v)$ where $v^{\prime}{}^{\frown}c$ is a lasso such that:
@@
The definition above is modeled on the statement of Theorem 5.9, which in turn is structured following the statement of (Rosendal’s criterion). (Theorem 5.6): the pairs $(\mathcal{U},u),(\mathcal{V},v)$ in the Definition correspond to their counterparts in the Theorem.
@@
The existence of a robust cycle gives us a brick partition $\mathcal{S}$, a spaced circular path $C$ of circular length $\ell\geq 4$, a brick partition $\mathcal{U}\preceq\mathcal{S}$, and and a walk $u$ on $\mathcal{U}$. Up to taking a refinement, by (Monotone Lifting). (Lemma 3.1) we can suppose that
@@
Suppose we are given a brick partition and a walk $(\mathcal{V},v_{0})\preceq(\mathcal{U},u)$. By definition of robust cycle, we can assume that, up to a refinement, $v_{0}$ is a lasso $v{}^{\frown}c_{+}$ such that:
@@
Let $c_{-}$ be the circular path going in the opposite direction of $c_{+}$. Set $N\coloneqq\operatorname{wd}_{C}(\rho^{\mathcal{V}}_{\mathcal{S}}\cdot c_{+})$, so that $\operatorname{wd}_{C}(\rho^{\mathcal{V}}_{\mathcal{S}}\cdot c_{-})=-N$. Up to switching the role of $c_{+},c_{-}$, we might assume that $N\geq\ell>0$.
@@
Since $v{}^{\frown}\bar{w}_{0}$ is uncrossed, by (Upgrading Walks to Paths). (Theorem 6.2) there is a brick partition $\widetilde{\mathcal{W}}\preceq\mathcal{V}$ and a spaced path $\tilde{w}_{0}$ on $\widetilde{\mathcal{W}}$ such that $\rho^{\widetilde{\mathcal{W}}}_{\mathcal{V}}\cdot\tilde{w}_{0}$ monotonically refines $v{}^{\frown}\bar{w}_{0}$. Since $v{}^{\frown}\bar{w}_{1}$ is uncrossed as well, by (Monotone Lifting). it can be monotonically refined by an uncrossed walk $\tilde{w}_{1}$ on $\widetilde{\mathcal{W}}$. Hence, applying (Upgrading Walks to Paths). to $\tilde{w}_{1}$, we can find a brick partition $\mathcal{W}\preceq\widetilde{\mathcal{W}}$ and a spaced path $w_{1}$ such that $\rho^{\mathcal{W}}_{\mathcal{V}}\cdot w_{1}$ monotonically refines $v{}^{\frown}\bar{w}_{1}$. Another application of (Monotone Lifting). to $\tilde{w}_{0}$ yields finally a spaced path $w_{0}$ on $\mathcal{W}$ such that $\rho^{\mathcal{W}}_{\mathcal{V}}\cdot w_{0}$ monotonically refines $v{}^{\frown}\bar{w}_{0}$. We stress that $w_{0},w_{1}$ are paths, not lassos.
@@
Sketch of Proof.
We only show that (b0) holds, the other cases being analogous. Let such a $t$ be given. By definition, $\operatorname{wd}_{C}(\rho^{\mathcal{V}}_{\mathcal{S}}\cdot\bar{w}_{0})>3M_{v}+M_{c}+2$, and for any final segment $s$ of $v$, $\operatorname{wd}_{C}(\rho^{\mathcal{V}}_{\mathcal{S}}\cdot s)\geq-M_{v}$. Since $\rho^{\mathcal{W}}_{\mathcal{V}}\cdot t$ monotonically refines a final segment of $v{}^{\frown}\bar{w}_{0}$ containing $\bar{w}_{0}$, we conclude by Lemma 3.3. ∎
We now apply Theorem 5.9; hence towards a contradiction suppose there are a brick partition $\mathcal{Z}$, two monotone epimorphisms $\alpha_{0},\alpha_{1}\colon\mathcal{Z}\to\mathcal{W}$, and a reduced walk $z$ on $\mathcal{Z}$ such that $\alpha_{i}\cdot z\sqsubseteq w_{i}$, for $i=0,1$, and
@@
Notice that $\bar{z}_{v}$ might be the empty walk (this is the case if $C=\mathcal{S}$). In particular, $\alpha_{0}\cdot\bar{z}_{0}$ refines and is confined to a final segment of $w_{0}$ whose image under $\rho^{\mathcal{W}}_{\mathcal{V}}$ contains $\bar{w}_{0}$, and moreover $\alpha_{0}\cdot\bar{z}_{0}(-1)=w_{0}(-1)$, so by 6.4.1.(b0) and Lemma 3.4
@@
$d_{\mathcal{U}}(\rho^{\mathcal{W}}_{\mathcal{U}}\alpha_{0}\cdot\bar{z}_{0}(j),\rho^{\mathcal{W}}_{\mathcal{U}}\alpha_{1}\cdot\bar{z}_{0}(j+1))\leq d_{\mathcal{U}}(\rho^{\mathcal{W}}_{\mathcal{U}}\alpha_{0}\cdot\bar{z}_{0}(j),\rho^{\mathcal{W}}_{\mathcal{U}}\alpha_{0}\cdot\bar{z}_{0}(j+1))+\\ +d_{\mathcal{U}}(\rho^{\mathcal{W}}_{\mathcal{U}}\alpha_{0}\cdot\bar{z}_{0}(j+1),\rho^{\mathcal{W}}_{\mathcal{U}}\alpha_{1}\cdot\bar{z}_{0}(j+1))\leq 2,$
@@
If $\bar{z}_{v}$ is the empty walk then $\operatorname{wd}_{C}(\rho^{\mathcal{W}}_{\mathcal{S}}\alpha_{1}\cdot\bar{z})=\operatorname{wd}_{C}(\rho^{\mathcal{W}}_{\mathcal{S}}\alpha_{1}\cdot\bar{z}_{0})$, so we obtain a contradiction from (6.6) and (6.9). Otherwise $\alpha_{1}\cdot\bar{z}_{v}$ refines an initial segment of $w_{1}$ and $\rho^{\mathcal{W}}_{\mathcal{V}}\alpha_{1}(\bar{z}_{v}(-1))$ is contained in $v$, and not in $\bar{w}_{1}$: indeed by definition of $\bar{z}_{0}$ the vertex $\rho^{\mathcal{W}}_{\mathcal{U}}(\alpha_{0}(\bar{z}_{v}(-1)))$ must be $E$-adjacent to one in $\mathcal{U}\setminus(\rho^{\mathcal{U}}_{\mathcal{S}})^{-1}(C)$, and this in combination with (6.5) yields
@@
Combining (6.9) and (6.10) we conclude $\operatorname{wd}_{C}(\rho^{\mathcal{W}}_{\mathcal{S}}\alpha_{1}\cdot\bar{z})>M_{v}+M_{c}$, again a contradiction with (6.6). ∎
@@ 6.3. Locally non-planar open subset
Useful examples of Peano continuum with no locally separating points which have locally non-planar open subsets to keep in mind throughout this subsection are the Menger curve, or any $n$-dimensional closed manifold, for $n\geq 3$.
@@
Proposition 6.5 (Disjoint Paths Property).
@@
For each $1\leq i<\operatorname{lh}(u)$, let $x_{i}$ be a point in $\operatorname{Cl}(u(i-1))\cap\operatorname{Cl}(u(i))\cap(u(i-1)\vee u(i))$, and, for each $1\leq j<\operatorname{lh}(u^{\prime})$, let $x^{\prime}_{j}$ be a point in $\operatorname{Cl}(u^{\prime}(j-1))\cap\operatorname{Cl}(u^{\prime}(j))\cap(u^{\prime}(j-1)\vee u^{\prime}(j))$. We can assume that all these points are distinct by 4.2.(2), since $u$ and $u^{\prime}$ are paths, so for each $i<\operatorname{lh}(u)$ there is at most one $j<\operatorname{lh}(u^{\prime})$ such that $u(i)=u^{\prime}(j)$, and vice versa.
@@
Let $\lambda,\lambda^{\prime}$ be the two disjoint arcs resulting from the concatenations of the $\gamma_{i}$’s and $\gamma^{\prime}_{j}$’s, respectively. Let $\tilde{\mathcal{V}}\preceq\mathcal{U}$ be a brick partition fine enough that $\operatorname{Star}(\lambda,\tilde{\mathcal{V}})\cap\operatorname{Star}(\lambda^{\prime},\tilde{\mathcal{V}})=\emptyset$. Set $P_{i}\coloneqq\operatorname{Star}(\lambda,\tilde{\mathcal{V}}(u(i))),P^{\prime}_{j}\coloneqq\operatorname{Star}(\lambda^{\prime},\tilde{\mathcal{V}}(u^{\prime}(j)))$, for each $i<\operatorname{lh}(u),j<\operatorname{lh}(u^{\prime})$, and let $\mathcal{V}$ be the amalgam of $\tilde{\mathcal{V}}$ refining $\mathcal{U}$ and containing $\bigvee P_{i}$ and $\bigvee P^{\prime}_{j}$, for all $i<\operatorname{lh}(u),j<\operatorname{lh}(u^{\prime})$. Then $v\coloneqq\langle\bigvee P_{0},\dots,\bigvee P_{\operatorname{lh}(u)-1}\rangle$ and $v^{\prime}\coloneqq\langle\bigvee P^{\prime}_{0},\dots,\bigvee P^{\prime}_{\operatorname{lh}(u^{\prime})-1}\rangle$ are paths on $\mathcal{V}$ satisfying the conclusions of the proposition.
@@
The (Disjoint Paths Property). allows us to find a robust cycle in any locally non-planar open set.
@@
Suppose that $O\subseteq X$ is open and locally non-planar. Fix any brick partition $\mathcal{W}$ of $X$ which is fine enough that there are $W,W^{\prime}$ which are in the same $E$-connected component of $\mathcal{W}(O)$ but are not adjacent. Let $w$ be a path from $W$ to $W^{\prime}$ in $\mathcal{W}(O)$. By (Path Doubling). (Proposition 6.1), there are a brick partition $\mathcal{S}\preceq\mathcal{W}$ and two paths $s_{0},s_{1}$ on $\mathcal{S}$ which are disjoint except for their endpoints and such that $\rho^{\mathcal{S}}_{\mathcal{W}}\cdot s_{i}$ monotonically refines $w$, for $i<2$. Let $C$ be the circular path which follows $s_{0}$ forward and then $s_{1}$ backwards. Up to a refinement, we can assume $C$ is spaced, by (Spacing out Paths). (Lemma 5.5). Since $s_{0},s_{0}$ have length at least $3$, $C$ has circular length at least $4$.
@@
Let $u$ be the empty walk and suppose that a brick partition and a walk $(\mathcal{V},v)\preceq(\mathcal{U},u)$ are given. We claim that it suffices to find $(\mathcal{V}^{\prime},v^{\prime})\preceq(\mathcal{V},v)$ with $v^{\prime}$ a path, and a circular path $c$ disjoint from $v^{\prime}$ satisfying (6.1) and (6.2) in Definition 6.3. Indeed, in this case, by (Path Doubling). we can assume that, up to a refinement, there are two paths $v^{\prime}_{0},v^{\prime}_{1}$ which are disjoint except in their endpoints, disjoint from $c$, and such that $\rho^{\mathcal{V}^{\prime}}_{\mathcal{V}}\cdot v^{\prime}_{i}\sqsubseteq v$. Let $P$ be a path from $v^{\prime}_{0}\cup v^{\prime}_{1}$ to $c$, and say without loss of generality that it starts in $v^{\prime}_{0}$. Then the lasso that follows $v^{\prime}_{1}$ from beginning to end, then $v^{\prime}_{0}$ backwards until meeting $P$, then follows $P$, then does a complete lap following $c$, is as required.
We now proceed to find $v^{\prime}$ and $c$ as above. Up to a refinement, we can suppose that $v$ is a path, by (Upgrading Walks to Paths). (Theorem 6.2). Use (Monotone Lifting). (Lemma 3.1) to find a spaced circular path $D^{\prime}$ on $\mathcal{V}$ such that $\rho^{\mathcal{V}}_{\mathcal{U}}\cdot D^{\prime}$ monotonically refines $D$. Since $D^{\prime}$ is confined in $\mathcal{V}(O)$, by (Disjoint Paths Property). (Proposition 6.5) there are a brick partition $\mathcal{V}^{\prime}\preceq\mathcal{V}$ and two disjoint paths $v^{\prime},c$ on $\mathcal{V}^{\prime}$ such that $\rho^{\mathcal{V}^{\prime}}_{\mathcal{V}}\cdot v^{\prime},\rho^{\mathcal{V}^{\prime}}_{\mathcal{V}}\cdot c$ monotonically refine $v$ and $D^{\prime}$, respectively, and $c$ is circular. Then $c$ satisfies (6.1) by (6.11), and it follows by Lemma 3.3 that it satisfies (6.2). ∎
6.4. Planar open set with non-locally-separating simple closed curve
The next two results capture the essential aspects behind the assumption of the existence of a non-locally-separating simple closed curve in Theorem 1.2. Good examples to keep in mind are the closed disk and the Sierpiński carpet.
We start with an auxiliary definition.